OpenStudy (anonymous):

integrate: 4/((1-x)(1+x)^2)

OpenStudy (anonymous):

\[\int\limits_{}^{} 4/(1-x)(1+x)^{2}\]

OpenStudy (experimentx):

make partial fractions.

OpenStudy (anonymous):

oh yeah!thank you! i never think of partial fraction :/ i'll work on it now

OpenStudy (anonymous):

can i take the constant out?

OpenStudy (anonymous):

\[4\int\limits_{}^{}1/(1-x)(x+1)^{2}\]

OpenStudy (experimentx):

should be easy http://www.mathsisfun.com/algebra/partial-fractions.html

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i found A=1 but got nothing for B and C

OpenStudy (anonymous):

i used linear system by say x=1, x=-1 and x=0

OpenStudy (experimentx):

Find A B and C \[ \frac{4}{((1-x)(1+x)^2)}\ = \frac{A}{(1-x)} + \frac{B}{(1+x)} + \frac{C}{(1+x)^2} \]